OpenStudy (anonymous):

Solve the system of equations for y. -3x-2y=-16 4x+3y=23

OpenStudy (anonymous):

which equation do you want to set equal to y?

OpenStudy (anonymous):

either one. I've been stuck on this problem for the longest!

OpenStudy (anonymous):

lets do the top one. -3x-2y=-16 -2y=-3x-16 y=3/2x+8

OpenStudy (anonymous):

now that you know what y is, plug it in 4x+3y=23

OpenStudy (anonymous):

wouldn't you be adding -2y to both sides? why is it still -2y?

hero (hero):

-3x-2y=-16 4x+3y=23 1st: Add both equations together to get: x + y = 7

hero (hero):

2nd - Isolate x: x = 7 - y 3rd - substitute the expression for x into one of the original equations 4th - Solve for y

OpenStudy (anonymous):

thank you both!

hero (hero):

Did you figure it out yet? What's the value of y?

OpenStudy (anonymous):

positive 5

hero (hero):

-3(7-y)-2y=-16 -21 + 3y - 2y = -16 -21 + y = -16 y = -16 + 21 y = 5 Correct

OpenStudy (anonymous):

thanks :)

hero (hero):

yw